OpenStudy (anonymous):

What's the derivative of (sin(x))^(57)?

Directrix (directrix):

57 [ sin(x) ] ^ (56) [ cos(x) ]

OpenStudy (anonymous):

that answer was wrong...

OpenStudy (anonymous):

chain rule: \[(\sin(x))^{57}=57(\sin(x))^{56}\times \cos(x)\]

Directrix (directrix):

Starlight says that answer is incorrect. I think it is correct - just a notational difference in writing (sin (x) ) ^ 57.

OpenStudy (anonymous):

\[g(x) = x ^{57}\]\[f(x)=\sin (x)\]\[g(f(x))=(\sin (x))^{57}\]\[g \prime(f(x)) = g \prime(f(x))f \prime(x)\]\[g \prime(f(x))=57(\sin(x))^{56}\times \cos(x)\]

OpenStudy (anonymous):

I don't know but it was cos(x) lol... http://www.wolframalpha.com/input/?i=57th+derivative+of+sin%28x%29 and thanks